12 15k 2ok 7find sum of forces and each direction. and moment of each axis. find static equilbrum.
Answer #1

Force in x direction=-20 kips=20 kips(acting in negative x direction)

Force in y direction=35 kips

Force in z direction=-15 kips=15 kips(acting in negative z direction)

Moment in member about x axis=0

Moment about y axis=20*8+15*10=310 kip-ft

Moment about z axis=35*12=420 kip-ft